# EARL: Entropy-Aware RL Alignment of LLMs for Reliable RTL Code Generation

### v1

@@ IV-A Overview of EARL
TABLE I: Comparative analysis of Verilog code generation performance. Bold indicates best result. Underline indicates the second best result.
@@
Supervised Fine Tuning VerilogEval [verilogeval] 16B 46.2 67.3 28.8 45.9 N/A N/A
ChipNeMo [chipnemo] 70B 53.8 N/A 27.6 N/A 28.0 20.7
RTLLLM [rtllm] 13B 65.3 77.2 43.7 51.8 52.6 40.7
RTLCoder [xie2024rtlcoder] 6.7B 37.2 64.9 16.9 35.7 89.1 51.7
OriGen [pku2024origen] 7B 43.1 55.1 30.8 31.9 96.5 51.7
VeriGen [nyu2024] 16B 44.0 52.6 30.3 43.9 N/A N/A
HaVen [haven2025] 6.7B 66.1 81.1 43.1 55.1 81.4 38.6
BetterV [2024betterv] 7B 68.1 79.4 46.1 53.7 N/A N/A
AutoVCoder [2024autovcoder] 7B 68.7 79.9 48.5 55.9 100 51.7
Reinforcement Learning VeriReason [rl2025verireason] 7B 69.8 83.1 47.9 58.4 N/A N/A
VeriSeek [2025rl_1] 6.7B 61.6 76.9 N/A N/A 94.8 54.2
@@
IV-B Entropy-Aware RL
Our entropy study in Section III shows that only a minority of high-entropy tokens drive effective learning. Standard RLVR pipelines update all tokens uniformly, which dilutes learning signals and wastes gradient budget on low-impact positions. To address this limitation, we introduce an entropy-based masking mechanism that selectively emphasizes high-entropy tokens during policy optimization. For each rollout response $o^{i}$, let $H_{t}^{i}$ denote the entropy of token $t$. We adopt response-level quantiles to mitigate cross-prompt heterogeneity. We compute a response-level quantile threshold $\tau_{\rho}^{i}$ as $\text{Quantile}({H_{t}^{i}}_{t=1}^{|o^{i}|},\rho)$, where $\rho\in(0,1)$ denotes the quantile level. Specifically, tokens whose entropy exceeds $\tau_{\rho}^{i}$ receive full gradient updates, while low-entropy tokens are down-weighted to zero, preserving their stable distributions learned by SFT. Formally, EARL extends a DAPO-style objective with entropy gating and KL regularization:
$J_{\mathrm{EARL}}(\theta)$ $=\mathbb{E}_{(q,a)\sim\mathcal{D},\{o^{i}\}_{i=1}^{G}\sim\pi_{\text{old}}(\cdot|q)}\Bigg[\frac{1}{\sum_{i=1}^{G}|o^{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o^{i}|}$
@@ IV-B Entropy-Aware RL
$-\beta\,\mathbb{D}_{\text{KL}}(\pi$ $\|\pi_{\text{ref}})\Big)\Bigg]$
$\text{s.t.}\quad 0<$ $\big|\{i\in\{1,\ldots,G\}\mid\texttt{is\_equivalent}(o^{i},a)\}\big|<G.$
@@
The indicator $\mathbb{I}[H_{t}^{i}\geq{\tau}_{\rho}^{i}]$ is the entropy mask whih differentiates the optimization strength across token type. The constraint enforces mixed-quality groups (both passing and failing samples) so that group-normalized advantages are meaningful and gradients do not degenerate.
@@
Note that this design is optimizer-agnostic. EARL can be applied to policy gradient objectives without modifying their core update rules. In practice, we instantiate EARL with both PPO and DAPO. By concentrating gradient updates on uncertain yet functionally critical tokens, EARL achieves sharper credit assignment, reduces wasted updates on deterministic syntax, and improves the stability of RLVR training for long, structured RTL sequences.
IV-C Reward Design
A central component of EARL is the design of verifiable rewards that align generation with downstream RTL correctness. Inspired by industrial EDA flows, we construct a cascaded reward hierarchy that mirrors the standard compilation–simulation–verification pipeline.
@@ IV-C Reward Design
Given a natural language specification $q$, the model generates a candidate Verilog program $o$. To obtain the reward of the program, we evaluate it through three stages: (1) Syntax Validity. The candidate is compiled using iverilog. Programs that fail syntax checks are immediately assigned zero reward, preventing wasted computation on invalid code. (2) Interface Consistency. If syntax passes, we check whether the module name and port declarations match the specification. Partial matches receive intermediate reward, while fully aligned interfaces receive higher credit. (3) Functional Equivalence. Finally, we invoke the eqy engine from Yosys to verify functional equivalence against a reference implementation. Passing candidates receive maximal reward, while near-misses are graded slightly lower to encourage exploration without reward hacking.
IV-D Dataset Preparation
@@
Training EARL requires functionally verified data to bootstrap supervised initialization and support reward evaluation during RL. To this end, we construct a synthetic dataset of specification–code–testbench triplets. The dataset generation pipeline follows two steps. Synthetic Generation. We employ GPT-4 to generate diverse prompts covering arithmetic operators, sequential logic, control modules, and algorithmic circuits. Each prompt is paired with a Verilog implementation and an executable testbench. Inspired by prior work [2024autovcoder, 2024betterv], we diversify prompts along multiple axes, including difficulty level, circuit type, and specification style. Compiler-Driven Filtering. To guarantee correctness, we apply a three-stage validation pipeline to all generated samples. First, syntax validity is verified using iverilog. Second, functional correctness is checked by executing the paired testbench. Finally, we employ a grading LLM to provide an additional layer of semantic evaluation on functional behavior. Only samples that pass all three stages are retained in the dataset.
V Experiments
V-A Experiment setup
Figure 5: Ablation study on RL algorithms and entropy mechanisms.
@@ V-A Experiment setup
Benchmarks. We evaluate on three standard RTL code generation benchmarks. VerilogEval v1 [verilogeval] includes 143 machine-generated problems (VerilogEval-Machine) and 156 expert-written tasks (VerilogEval-Human), designed to test structural and semantic generation capabilities. RTLLM v1.1 [rtllm] contains 29 diverse RTL design prompts focused on real-world functionality.
Metrics. Following prior works [chipnemo, rtllm], we use the pass@k metric to evaluate the probability of generating a correct solution within $k$ attempts, where $k\in\{1,5\}$. For each prompt, we generate $n=5$ completions and compute:
@@
V-B Comparison with Competing Methods
We evaluate the syntactic and functional correctness of EARL against a wide range of baseline and state-of-the-art methods, including supervised fine-tuning (SFT) approaches and reinforcement learning with verifiable rewards (RLVR). Note that for completeness, we also report results of prompt-engineering based methods, e.g., Origen, HaVen. Their pass-k metrics are obtained from our own evaluation under the same experimental setup. Results in Table I demonstrate that EARL achieves the best performance across all benchmarks and metrics. On VerilogEval-Machine, EARL attains 72.9% pass@1 and 83.9% pass@5, surpassing the strongest RL baseline VeriReason by 3.1% and 0.8%, respectively, and outperforming the best SFT model by 4.2% in pass@1. On VerilogEval-Human, which contains expert-written tasks more aligned with real-world HDL practices, EARL achieves 49.6% pass@1 and 60.2% pass@5, both the highest among all methods. On RTLLM v1.1, EARL obtains perfect syntax pass@5 (100%) and the highest functional correctness (68.9%), outperforming the best RL method by 14.7% and the best SFT method by 17.2%. These consistent improvements demonstrate that EARL establishes a new state-of-the-art in Verilog code generation, achieving reliable one-pass correctness while remaining parameter- and data-efficient.
TABLE II: Comparative analysis of threshold on Verilog code generation performance. Bold indicates best result. Underline indicates the second best result.
@@ V-B Comparison with Competing Methods
V-C Ablation Study
V-C1 Effect of RL Algorithms and Entropy Mechanisms
To better understand the effect of our entropy-aware reinforcement learning framework, we conduct an ablation study comparing different RL algorithms (PPO and DAPO) and different entropy mechanisms (Archer [2025stabilizing] and the proposed EARL). The results are summarized in Fig. 5.
@@ V-C1 Effect of RL Algorithms and Entropy Mechanisms
V-C2 Entropy Quantile
We further ablate the effect of the entropy quantile $\rho$ used to construct the response-level mask. Results in Table II show that performance peaks around $\rho=0.8$. Lower thresholds (e.g., 0.2) admit excessive noise, while higher thresholds (e.g., 1.0) exclude too many decision-critical tokens. This validates that selectively emphasizing a minority of high-entropy tokens yields the best trade-off between exploration and stability.
VI Conclusion

### v2

@@ 4.1. Overview of EARL
The following subsections describe each component of the framework in detail, including the formulation of entropy-aware RL, the design of verifiable rewards, and the dataset construction pipeline.
Table 1. Comparative analysis of Verilog code generation performance. Bold indicates best result. Underline indicates the second best result.
@@
Supervised Fine Tuning VerilogEval (Pinckney et al., 2025) 16B 46.2 67.3 28.8 45.9 N/A N/A
ChipNeMo (Liu et al., 2024b) 70B 53.8 N/A 27.6 N/A 28.0 20.7
RTLLLM (Lu et al., 2024) 13B 65.3 77.2 43.7 51.8 52.6 40.7
RTLCoder (Liu et al., 2024c) 6.7B 37.2 64.9 16.9 35.7 89.1 51.7
OriGen (Cui et al., 2025) 7B 43.1 55.1 30.8 31.9 96.5 51.7
VeriGen (Thakur et al., 2024) 16B 44.0 52.6 30.3 43.9 N/A N/A
HaVen (Yang et al., 2025) 6.7B 66.1 81.1 43.1 55.1 81.4 38.6
BetterV (Pei et al., 2024) 7B 68.1 79.4 46.1 53.7 N/A N/A
AutoVCoder (Gao et al., 2024) 7B 68.7 79.9 48.5 55.9 100 51.7
Reinforcement Learning VeriReason (Wang et al., 2025e) 7B 69.8 83.1 47.9 58.4 N/A N/A
VeriSeek (Wang et al., 2025c) 6.7B 61.6 76.9 N/A N/A 94.8 54.2
@@
4.2. Entropy-Aware RL
Our entropy study in Section 3 shows that only a minority of high-entropy tokens drive effective learning. However, existing RLVR pipelines apply uniform gradient updates across all tokens, which dilutes learning signals and wastes gradient budget on low-impact positions. To address this limitation, we introduce an entropy-gated policy optimization method that selectively emphasizes high-uncertainty tokens during policy optimization (see green block in Figure 4). For each rollout $\tilde{o}^{i}=(o_{1}^{i},\dots,o_{T_{i}}^{i})$, we compute the per-token entropy $H_{t}^{i}$ as defined in Eq. (4). To mitigate cross-prompt entropy scale variation, we adopt response-level quantile masking: a token $t$ is updated only if $H_{t}^{i}\geq\tau_{\rho}^{i}$, where $\tau_{\rho}^{i}=\operatorname{Quantile}(\{H_{t}^{i}\}_{t=1}^{T_{i}},\rho)$ for a chosen threshold $\rho\in(0,1)$. Specifically, tokens whose entropy exceeds $\tau_{\rho}^{i}$ receive full gradient updates, while low-entropy tokens are down-weighted to zero, preserving their stable distributions learned by SFT. Formally, EARL extends a DAPO-style objective with entropy gating :
(5) $J_{\mathrm{EARL}}(\theta)$ $=\mathbb{E}_{(q,a)\sim\mathcal{D},\{\tilde{o}^{i}\}_{i=1}^{G}\sim\pi_{\text{old}}(\cdot|q)}\Bigg[\frac{1}{\sum_{i=1}^{G}|\tilde{o}^{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|\tilde{o}^{i}|}$
@@ 4.2. Entropy-Aware RL
$-\beta\,\mathbb{D}_{\text{KL}}(\pi$ $\|\pi_{\text{ref}})\Big)\Bigg],\quad\text{s.t.}\quad 0<\left|\left\{i\mid R^{i}>0\right\}\right|<G,$
@@
where, the token-level advantage $\hat{A}_{t}^{i}$ follows Eq. (3). The indicator $\mathbb{I}[H_{t}^{i}\geq{\tau}_{\rho}^{i}]$ is the entropy mask which differentiates the optimization strength across token type. The constraint enforces diversity within each group, preventing degenerate all-pass or all-fail groups, which would invalidate the group-normalized advantage.
@@
Note that this design is optimizer-agnostic. EARL can be applied to policy gradient objectives without modifying their core update rules. In our experiments, we instantiate EARL with both PPO and DAPO. By concentrating gradient updates on uncertain yet functionally critical tokens, EARL achieves sharper credit assignment, reduces wasted updates on deterministic syntax, and improves the stability of RLVR training for long, structured RTL sequences.
4.3. Reward Design
@@
A central component of EARL is the design of verifiable rewards that align generation with downstream RTL correctness. Inspired by industrial EDA flows, we construct a cascaded reward hierarchy (see red block of Figure 4), which mirrors the standard compilation–simulation–verification pipeline. Given a natural language specification $q$, the model generates a rollout $\tilde{o}$ sampled from its current policy. To assign a reward, we evaluate $\tilde{o}$ through a cascaded three-stage verification process: First, we check syntax validity by compiling the candidate with iverilog. Programs that fail syntax checks are immediately assigned zero reward, preventing wasted computation on invalid code. If syntax passes, we proceed to evaluate interface consistency. We examine whether the module name and port declarations match the specification. Partial matches receive intermediate reward, while fully aligned interfaces receive higher credit. Finally, we assess functional equivalence. We invoke the eqy engine from Yosys to compare the output against a reference implementation. Passing candidates receive maximal reward, while near-misses are graded slightly lower to encourage exploration without reward hacking.
4.4. Dataset Preparation
@@
Training EARL requires functionally verified data to bootstrap supervised initialization and support reward evaluation during RL. To this end, we construct a synthetic dataset of specification–code– testbench triplets (see yellow block of Figure 4). The dataset generation pipeline follows two steps. The first step is synthetic generation. We prompt GPT-4 to generate diverse problem descriptions covering arithmetic operators, sequential logic, control modules, and algorithmic circuits. Each problem description is paired with a Verilog implementation and an executable testbench. Inspired by prior work (Gao et al., 2024; Pei et al., 2024), we diversify the prompts along multiple axes, including difficulty level, circuit type, and specification style. The second step involves compiler-driven filtering. To guarantee correctness, we apply a three-stage validation pipeline to all generated samples. First, syntax validity is verified using iverilog. Second, functional correctness is checked by executing the paired testbench. Finally, we employ a grading LLM to provide an additional layer of semantic evaluation on functional behavior. Only samples that pass all three stages are retained in the dataset.
5. Experiments
5.1. Experiment setup
Figure 5. Ablation study on RL algorithms and entropy mechanisms.
@@ 5.1. Experiment setup
Benchmarks. We evaluate on three standard RTL code generation benchmarks. VerilogEval v1 (Pinckney et al., 2025) includes 143 machine-generated problems (VerilogEval-Machine) and 156 expert-written tasks (VerilogEval-Human), designed to test structural and semantic generation capabilities. RTLLM v1.1 (Lu et al., 2024) contains 29 diverse RTL design prompts focused on real-world functionality.
Metrics. Following prior works (Liu et al., 2024b; Lu et al., 2024), we use the pass@k metric to evaluate the probability of generating a correct solution within $k$ attempts, where $k\in\{1,5\}$. For each prompt, we generate $n=5$ completions and compute:
@@
5.2. Comparison with Competing Methods
We evaluate the syntactic and functional correctness of EARL against a wide range of baseline and state-of-the-art methods, including SFT and RLVR approaches. Note that for completeness, we also report results of prompt-engineering based methods, e.g., Origen, HaVen. Their pass-k metrics are obtained from our own evaluation under the same experimental setup. Results in Table 1 demonstrate that EARL achieves the best performance across all benchmarks and metrics. On VerilogEval-Machine, EARL attains 72.9% pass@1 and 83.9% pass@5, surpassing the strongest RL baseline VeriReason by 3.1% and 0.8%, respectively, and outperforming the best SFT model by 4.2% in pass@1. On VerilogEval-Human, which contains expert-written tasks more aligned with real-world HDL practices, EARL achieves 49.6% pass@1 and 60.2% pass@5, both the highest among all methods. On RTLLM v1.1, EARL obtains perfect syntax pass@5 (100%) and the highest functional correctness (68.9%), outperforming the best RL method by 14.7% and the best SFT method by 17.2%. These consistent improvements demonstrate that EARL establishes a new state-of-the-art in Verilog code generation, achieving reliable one-pass correctness while remaining parameter- and data-efficient.
Table 2. Comparative analysis of threshold on Verilog code generation performance. Bold indicates best result. Underline indicates the second best result.
@@ 5.2. Comparison with Competing Methods
5.3. Ablation Study
5.3.1. Effect of RL Algorithms and Entropy Mechanisms.
To better understand the effect of our entropy-aware reinforcement learning framework, we conduct an ablation study comparing different RL algorithms (PPO and DAPO) and different entropy mechanisms (Archer (Wang et al., 2025a) and the proposed EARL). The results are summarized in Fig. 5.
@@ 5.3. Ablation Study
5.3.2. Entropy Quantile.
We further ablate the effect of the entropy quantile $\rho$ used to construct the response-level mask. Results in Table 2 show that performance peaks around $\rho=0.8$. Lower thresholds (e.g., 0.2) admit excessive noise, while higher thresholds (e.g., 1.0) exclude too many decision-critical tokens. This validates that selectively emphasizing a minority of high-entropy tokens yields the best trade-off between exploration and stability.
6. Conclusion